Is the relationship between te varibables in the table a direct variation an inverse or neither? if it is a direct or inverse write a functon to model it.
x= -9 -5 0 3
Y= -3 (-5/3) 0 1
(-5/3) is a fraction

Answer by stanbon(48558)

Is the relationship between te varibables in the table a direct variation an inverse or neither? if it is a direct or inverse write a functon to model it.
x= -9 -5 0 3
Y= -3 (-5/3) 0 1
-------------
y=x/3
or
y=(1/3)x
x and y have a direct relation.
That means "as x increases, y increases".
Cheers,
Stan H.